OpenStudy (anonymous):

What is t equal to if t+10-4t= -11

OpenStudy (anonymous):

Add t + -4t = -3t -3t +10 = -11 -3t = -21 t=7

OpenStudy (anonymous):

thnk! :D

OpenStudy (anonymous):

Thankyou